# Pressure-temperature phase diagram calculations using polynomial machine learning potentials: A comprehensive study based on global structure prediction and self-consistent phonon calculations

### v1

@@ IV.4 Accuracy of SSCHA calculations
IV.4.1 Required number of sample structures
@@
IV.4.2 Variance of free energy
Due to the structure sampling approach and the convergence tolerance parameter employed in the SSCHA, the calculated free energy exhibits fluctuations that depend on the specific simulation run. Figure 14(b) presents the standard deviations of the free energy values across the local minimum structures that exhibit dynamically stable states, along with the number of FC basis vectors. The standard deviations are derived from free energy values obtained from 10 independent SSCHA trials started with random FC values. In each trial, the number of sample structures per iteration is set between 5,000 and 20,000, as used in the current SSCHA calculations. All standard deviation values are below 1 meV/atom, indicating that the current SSCHA calculations yield robust free energy values. Note that these free energy values are utilized for EOS fittings, and the fitted free energy values are subsequently used to evaluate phase stability. Consequently, errors arising from free energy fluctuations are expected to be minimized through the EOS fitting process.
IV.4.3 Predictive power
Figure 15: Error distributions for energy and force values of sample structures in SSCHA calculations for local minimum structures. The energy and force values are computed using (a) the general-purpose MLP and (b) the specialized MLPs, and compared to those obtained from DFT calculations. The numerical values enclosed in the squares indicate RMSEs for energy and force, expressed in meV/atom and $\rm{\AA}$, respectively.
We evaluate the accuracy of both general-purpose and specialized MLPs for SSCHA calculations. In the SSCHA, energy and force values from sample structures are utilized to determine the free energy and effective FCs. Figure 15 illustrates the error distributions for energy and force values from representative sample structures obtained in SSCHA calculations for local minimum structures. For each combination of 81 local minimum structures and 10 to 20 volume conditions, one representative sample structure is selected from those generated by the converged FCs at 1000 K. In Fig. 15(a), energy and force values computed using the general-purpose MLP are compared with those from DFT calculations. Despite many local minimum structures differing from the prototype structures used to develop the general-purpose MLP, the MLP demonstrates reasonable predictive capability for SSCHA calculations. This indicates that the general-purpose MLP can enable reliable SSCHA calculations across a wide range of local minimum structures. However, prediction errors for energy values tend to be slightly larger for structures with low cohesive energies, which correspond to structures with volumes near equilibrium at low pressure. Figure 15(b) illustrates the performance of the specialized MLPs in predicting the energies and forces of representative sample structures in 26 local minimum structures. As shown in Fig. 15(b), the predictive accuracy of the specialized MLPs is enhanced relative to the general-purpose MLP, thereby enabling more accurate SSCHA calculations.
We also examine the effect of DFT dataset size on prediction errors in constructing the specialized MLPs. Figure 16 shows the relationship between the RMSE evaluated by the cross-validation approach and the DFT dataset size for five local minimum structures. The dataset sizes are expressed as multiples of the current DFT dataset size used to develop the specialized MLPs, ranging from 0.2 to 2.0. The predictive accuracy has sufficiently converged with the current dataset size. Additionally, MLPs developed with dataset sizes as small as 0.4 exhibit reasonable predictive performance, indicating that even smaller datasets can yield sufficiently accurate predictions.
@@ V Conclusion
This study has demonstrated a procedure for performing the efficient and accurate enumeration of crystal structures using polynomial MLPs over a wide range of pressures. Reliable polynomial MLPs, capable of comprehensively enumerating structures across a wide pressure range, have been developed using an iteratively constructed DFT dataset. The dataset includes random structures derived from optimized prototype structures at pressures ranging from 0 to 100 GPa, as well as local minimum structures obtained from RSS. To flexibly represent interatomic interactions in diverse structures, we have proposed hybrid polynomial MLP models that integrate multiple polynomial MLP models. The MLP has successfully identified a variety of globally stable and metastable structures, including almost all experimentally reported structures, with accuracy comparable to DFT calculations.
This study has also introduced a procedure that integrates polynomial MLP to achieve phase stability analysis through global structure search. We have demonstrated the application of this method to evaluate the phase stability of elemental silicon across a temperature range of 0–1000 K and a pressure range of 0–100 GPa, using local minimum structures obtained from the global structure search. In the current approach, the SSCHA method, combined with polynomial MLP, is employed to systematically assess the free energy values of these local minimum structures. The SSCHA method accounts for anharmonic vibrational contributions and enables the evaluation of free energies for dynamically stabilized structures by considering the temperature dependence of effective phonon frequencies. A comprehensive free-energy evaluation of the local minimum structures has yielded a reliable pressure-temperature phase diagram.
The polynomial MLP is crucial to the current approach. This approach involves global structure searches with numerous local geometry optimizations and systematic SSCHA calculations for many structures under various conditions. Additionally, each SSCHA calculation requires the sampling of a large number of supercell structures. Since all these procedures demand extensive energy and force evaluations, the robust polynomial MLPs employed in this study have proven highly efficient in facilitating these processes while maintaining high accuracy. Consequently, this study provides a general and robust framework for reliable crystal structure predictions under pressure and finite-temperature conditions, applicable to a wide range of systems.
@@ Acknowledgements.
Appendix A Leave-one-out cross validation
Here, we introduce the weighted leave-one-out CV method to estimate the prediction error of the polynomial MLP. The weighted leave-one-out CV score is defined by
(CV $(\text{CV score})=\frac{1}{N_{\rm{data}}}\sum_{i=1}^{N_{\rm{data}}}\left\{w_{i}(\bm{x}_{i}^{T}\hat{\bm{\omega}}_{(-i)}-y_{i})\right\}^{2},$ (16)
where $N_{\text{data}}$ represents the number of the training data points. $\bm{x}_{i}$, $y_{i}$, and $w_{i}$ correspond to the predictor vector, the observation value, and the weight for the $i$-th data point, respectively. The vector $\hat{\bm{\omega}}_{(-i)}$ denotes the regression coefficients estimated using the dataset excluding the $i$-th data point.

### v2

@@ IV.4 Accuracy of SSCHA calculations
IV.4.1 Predictive power
Figure 8: Distributions of (a) energies and (b) forces for sample structures in SSCHA calculations for local minimum structures. The energy and force values are computed using various interatomic potentials, including the general-purpose polynomial MLP and the specialized polynomial MLPs, and are compared with those obtained from DFT calculations. These potentials include the MEAM potential [43], the Tersoff potential [41], the quadratic SNAP [44], and the GAP [18]. The numerical values enclosed in the squares indicate RMSEs for energy and force, expressed in meV/atom and $\rm{\AA}$, respectively. In the figure, “GP-pMLP”, “S-pMLP”, and “Quad. SNAP” stand for general-purpose polynomial MLP, specialized polynomial MLP, and quadratic SNAP, respectively.
We evaluate the accuracy of both general-purpose and specialized MLPs in SSCHA calculations. In the SSCHA, energy and force values of sample structures are utilized to determine the free energy and effective FCs. Figure 8 illustrates the distributions of (a) energies and (b) forces for representative sample structures obtained from SSCHA calculations for local minimum structures, computed using the general-purpose MLP, the specialized MLPs, and DFT calculations. In the panel illustrating the distribution obtained from the general-purpose MLP, one representative sample structure is selected for each combination of 81 local minimum structures and 10 to 20 volume conditions, from those generated using the converged FCs at 1000 K. Despite many local minimum structures differing from the prototype structures used to develop the general-purpose MLP, the MLP demonstrates reasonable predictive capability for SSCHA calculations. This indicates that the general-purpose MLP can enable reliable SSCHA calculations across a wide range of local minimum structures. However, as shown in the supplemental material, prediction errors for energy values tend to be slightly larger for structures with low cohesive energies, which correspond to structures with volumes near equilibrium at low pressure.
To evaluate the predictive accuracy of the specialized MLPs in SSCHA calculations, representative sample structures obtained from 26 local minimum structures are employed. The predictive accuracy of the specialized MLPs is enhanced relative to that of the general-purpose MLP, thereby enabling more accurate SSCHA calculations. The effect of the DFT dataset size on the prediction errors in constructing the specialized MLPs for five local minimum structures is presented in Appendix D, demonstrating that the predictive accuracy has sufficiently converged with the current dataset size.
We also evaluate the accuracy of interatomic potentials other than the polynomial MLP in SSCHA calculations. These include the MEAM potential [43], the Tersoff potential [41], the quadratic SNAP [44], and the GAP [18]. Figure 8 also shows the distributions of (a) energies and (b) forces for the same sample structures as those used in the panel illustrating the distribution obtained from the specialized polynomial MLPs, computed using these interatomic potentials and DFT calculations. All of these potentials fail to accurately predict the energies of the sample structures, particularly under high-pressure conditions. Furthermore, the forces predicted by these potentials fail to reproduce those obtained from DFT calculations.
IV.4.2 Required number of sample structures
@@
IV.4.3 Variance of free energy
Due to the structure sampling approach and the convergence tolerance parameter employed in the SSCHA, the calculated free energy exhibits fluctuations that depend on the specific simulation run. Figure 9(b) presents the standard deviations of the free energy values across the local minimum structures that exhibit dynamically stable states, along with the number of FC basis vectors. The standard deviations are derived from free energy values obtained from 10 independent SSCHA trials started with random FC values. In each trial, the number of sample structures per iteration is set between 5,000 and 20,000, as used in the current SSCHA calculations. All standard deviation values are below 1 meV/atom, indicating that the current SSCHA calculations yield reasonable free energy values. Note that these free energy values are utilized for EOS fittings, and the fitted free energy values are subsequently used to evaluate phase stability. Consequently, errors arising from free energy fluctuations are expected to be minimized through the EOS fitting process.
@@ V Conclusion
This study proposes a procedure for accurately assessing phase stability through global structure searches under various pressure and temperature conditions. The approach combines global structure searches involving numerous local geometry optimizations with systematic SSCHA calculations performed on many structures across different conditions. Central to this procedure is the development of reliable polynomial MLPs capable of enumerating diverse local minimum structures over a wide pressure range and enabling SSCHA calculations under combined pressure and temperature conditions. The developed MLPs significantly accelerate the extensive energy and force evaluations required, while maintaining high accuracy.
The performance of the proposed procedure has been demonstrated for elemental silicon as a prototypical example. A polynomial MLP was developed using a dataset comprising randomly generated structures derived from optimized prototype structures at pressures ranging from 0 to 100 GPa, in addition to local minimum structures obtained from RSS. The MLP successfully identified a wide variety of globally stable and metastable structures, including nearly all experimentally reported phases, with accuracy comparable to that of DFT calculations. Furthermore, systematic free-energy evaluations of the local minimum structures, incorporating anharmonic contributions, yielded a reliable pressure-temperature phase diagram spanning 0–1000 K and 0–100 GPa. Consequently, this study establishes a general and robust framework for reliable crystal structure prediction under pressure and finite-temperature conditions, applicable to a broad range of materials systems.
@@ Acknowledgements.
Appendix A Pareto-optimal MLPs
Figure 10: Prediction error and computational efficiency of polynomial MLPs obtained from the grid search in elemental Si. The blue and green open circles represent the properties of MLPs constructed using single polynomial MLP models and hybrid polynomial MLP models, respectively. Computational efficiency is assessed based on the computation time using a single core of the ${}^{\text{\textregistered}}$ ${}^{\text{\textregistered}}$ E5-2630 v3 (2.40 GHz). The orange open squares denote the Pareto-optimal MLPs across both single and hybrid polynomial MLP models, while the blue open diamonds represent the Pareto-optimal MLPs from the single polynomial MLP models only. The red filled square indicates the selected MLP for conducting the global structure search and phase stability analysis.
Figure 10 shows the prediction error and computational efficiency of the polynomial MLPs obtained through a grid search of 523 single-polynomial and 640 hybrid-polynomial MLP models for elemental Si. These MLPs are constructed using dataset 2, which includes approximately 16,000 structures shown in Fig. 2. The prediction errors are evaluated using the RMSEs of the energy. The computational efficiency is assessed by measuring the elapsed time to compute the energy, forces, and stress tensors ten times for a structure containing 1,000 atoms. As the model complexity of the polynomial MLP increases, the prediction error generally decreases. Among the Pareto-optimal MLPs, the polynomial MLPs constructed using hybrid polynomial MLP models demonstrate lower prediction errors than those constructed with single polynomial MLP models when the computational cost exceeds 1.5 ms/step/atom. Hence, the hybrid polynomial MLPs can provide Pareto-optimal MLPs with a better trade-off between prediction error and computational cost. The optimal MLP selected for global structure searches and phase stability analyses is indicated by the red filled square in Fig. 10.
Appendix B Predictive power of polynomial MLPs
Figure 12 presents the phonon density of states (DOS) for diamond-, $\beta$-Sn-, SH-, Si-VI-, HCP-, and FCC-type structures at 50 GPa in elemental Si. These calculations are performed using the DFT and the two MLPs within the harmonic approximation. The finite displacement method in the phonopy code [99] is employed to obtain phonon properties and thermal expansion, with supercells constructed by expanding the conventional unit cells of the respective structures. The phonon DOS calculated with the MLP constructed from dataset 1 is inconsistent with the DFT results, especially for the FCC-type structure. By comparison, the phonon DOS obtained using the MLP from dataset 2 agrees with the DFT results.
Figure 13 shows the temperature dependence of thermal expansion, calculated using the quasi-harmonic approximation, for the $\beta$-Sn-, SH-, HCP-, and FCC-type structures at 10, 30, 60, and 90 GPa, respectively. Although developing MLPs that accurately predict thermal expansion is more challenging than those designed to predict phonon DOS, the MLP developed from dataset 2 accurately predicts thermal expansion. By contrast, the MLP trained on dataset 1 produces inconsistent results and cannot compute the thermal expansion for the FCC-type structure because it predicts the FCC-type structure to be dynamically unstable.
Appendix C Leave-one-out cross validation
We introduce the weighted leave-one-out CV method to estimate the prediction error of the polynomial MLP. The weighted leave-one-out CV score is defined by
(CV $(\text{CV score})=\frac{1}{N_{\rm{data}}}\sum_{i=1}^{N_{\rm{data}}}\left\{w_{i}(\bm{x}_{i}^{T}\hat{\bm{w}}_{(-i)}-y_{i})\right\}^{2},$ (17)
where $N_{\text{data}}$ represents the number of the training data points. $\bm{x}_{i}$, $y_{i}$, and $w_{i}$ correspond to the predictor vector, the observation value, and the weight for the $i$-th data point, respectively. The vector $\hat{\bm{w}}_{(-i)}$ denotes the regression coefficients estimated using the dataset excluding the $i$-th data point.
@@ Appendix C Leave-one-out cross validation
Appendix D Effect of DFT dataset size on specialized MLP accuracy
The effect of DFT dataset size on the prediction errors in constructing the specialized MLPs is investigated. Figure 14 shows the relationship between the RMSE evaluated by the cross-validation approach and the DFT dataset size for five local minimum structures. The dataset sizes are expressed as multiples of the current DFT dataset size used to develop the specialized MLPs, ranging from 0.2 to 2.0. The predictive accuracy has sufficiently converged with the current dataset size. Additionally, MLPs developed with dataset sizes as small as 0.4 exhibit reasonable predictive performance, indicating that even smaller datasets can yield sufficiently accurate predictions.
Supplemental Material
Global structure searches under varying temperatures and pressures using polynomial machine learning potentials: A case study on silicon
D.1 Model parameters of the selected polynomial MLP
We select a hybrid polynomial MLP with a computational cost of 3.1 ms/atom/step from the set of Pareto-optimal MLPs, balancing prediction accuracy and efficiency, as discussed in Section II E of the main text. The root mean square errors (RMSEs) for energy, force, and stress tensor predictions are 2.8 meV/atom, 0.055 eV/Å, and 35.9 meV/atom, respectively. This hybrid polynomial model consists of two sub-models. One uses a cutoff radius of 5.0 ${\rm{\AA}}$, and the other uses 6.0 ${\rm{\AA}}$. Both use seven radial basis functions and spherical harmonics truncated at $l^{(2)}_{\rm{max}}=8$ and $l^{(3)}_{\rm{max}}=4$. The potential energy in each case is given by
$E^{(i)}=F_{1}\left(D^{(i)}\right)+F_{2}\left(D^{(i)}\right),$ (S21)
where the total number of expansion coefficients is 18,720. This model is designated as optimal for the present global structure searches and phase stability analyses.
D.2 Metastable structures
Figure S15 shows the relative enthalpy values from DFT calculations for local minimum structures shown in Fig. 5(b) of the main text. Structures whose relative enthalpy values exceed 10 meV/atom across the entire pressure range are excluded from Fig. S15. In addition to the stable structures listed in Table I of the main text, three metastable local minimum structures are identified in the 0–80 GPa range in Fig. S15(a). They correspond to the Si-IV structure, a structure with space group $I4_{1}/acd$ represented with 32 atoms, and a structure with space group $R\bar{3}m$. Moreover, it is also observed that the enthalpy values of the Si-XI structure are nearly identical to those of the $\beta$-Sn-type structures. Between 80 and 100 GPa in Fig. S15(b), 15 metastable structures with low relative enthalpy values are found. These structures exhibit different stacking orders of close-packed planes, indicating that elemental Si tends to adopt close-packed structures under high pressures.
Figure S16 illustrates the relationship between pressure and the number of local minimum structures with relative Gibbs free energies below 5 meV/atom, excluding the global minimum structure, as calculated using the SSCHA with specialized MLPs. The number of these local minimum structures remains nearly constant as the temperature increases. Additionally, many structures exhibit low relative Gibbs free energies in the 80–100 GPa range, where close-packed structures such as the HCP-, $\alpha$-La-, and FCC-type structures are stable. The maximum number of local minimum structures with relative Gibbs free energies below 5 meV/atom is 17, observed at 86.5 GPa and 900 K. Among these structures, 16 correspond to those with low relative enthalpy, as shown in Fig. S15, in the pressure range of 80–100 GPa at 0 K.
Figure S15: Relative enthalpy values computed using DFT calculations for local minimum structures. Local minimum structures with relative enthalpy values exceeding 10 meV/atom across the entire pressure range are excluded from the analysis. The legend indicates the space groups of the structures or their corresponding prototype structures, and $Z$ denotes the number of atoms in the unit cell. (a) Enthalpy values are shown relative to the Si-V structure for pressures from 0 to 80 GPa in the left panel. The relative enthalpies in the central panel are displayed for the 9–12 GPa range, where multiple structures have very similar enthalpy values. (b) Enthalpy values are shown relative to the $\alpha$-La-type structure for pressures from 80 to 100 GPa.
Figure S16: Relationship between pressure and the number of local minimum structures in elemental Si. Only structures with relative Gibbs free energies below 5 meV/atom are included, excluding the global minimum structure. The pressure ranges from 0 to 100 GPa, and the numbers are shown for temperatures of 200, 500, and 1000 K.
D.3 Energy errors of polynomial MLPs in SSCHA calculations
Figure S17 illustrates the distributions of energy errors for representative sample structures obtained from SSCHA calculations of local minimum structures. Energy values are computed using the general-purpose MLP and the specialized MLP, and are compared against those from DFT calculations. Although the general-purpose MLP demonstrates reasonable predictive performance in SSCHA calculations, its energy prediction errors tend to be slightly larger for structures with low cohesive energies, which correspond to volumes near equilibrium under low pressure. In contrast, the specialized MLPs accurately reproduce the DFT energy values across all local minimum structures, thereby enabling more accurate SSCHA calculations.
Figure S17: Distributions of energies for sample structures in SSCHA calculations for local minimum structures. The energy and force values are computed using the general-purpose polynomial MLP and the specialized polynomial MLPs, and are compared with those obtained from DFT calculations. The numerical values enclosed in the squares indicate RMSEs for energy, expressed in meV/atom. In the figure, GP-pMLP and S-pMLP stand for general-purpose polynomial MLP and specialized polynomial MLP, respectively.